# ProReason: Multi-Modal Proactive Reasoning with Decoupled Eyesight and Wisdom

### v3

@@ Abstract
Large vision-language models (LVLMs) have witnessed significant progress on visual understanding tasks. However, they often prioritize language knowledge over image information on visual reasoning tasks, incurring performance degradation. To tackle this issue, we first identify the drawbacks of existing solutions (i.e., limited multi-modal reasoning capacities, and insufficient and irrelevant visual descriptions). We then decompose visual reasoning process into two stages: proactive visual perception (i.e., eyesight) and textual reasoning (i.e., wisdom), and introduce a novel visual reasoning framework named ProReason. This framework features decoupled vision-reasoning capabilities and multi-run proactive perception. Briefly, given a multi-modal question, ProReason iterates proactive information collection and reasoning until the answer can be concluded with necessary and sufficient visual descriptions. Notably, the disassociation of capabilities allows seamless integration of existing large language models (LLMs) to compensate for the reasoning deficits of LVLMs. Our extensive experiments demonstrate that ProReason outperforms existing multi-step reasoning frameworks on various benchmarks for both open-source and closed-source models, with the average performance gain reaching 13.2%. Besides, the integration of LLMs allows ProReason to produce high-quality visual reasoning data, which empowers ProReason-distilled models (i.e., ProReason-VL and ProReason-Q3) to achieve superior performance in downstream tasks. Our insights into existing solutions and the decoupled perspective for feasible integration of LLMs illuminate future research on visual reasoning techniques, especially LLM-assisted ones.
@@
Jingqi Zhou, Sheng Wang, Jingwei Dong, Kai Lui, Lei Li, Jiahui Gao, Jiyue Jiang, Lingpeng Kong, Chuan Wu The University of Hong Kong, The Chinese University of Hong Kong, Xi’an Jiaotong University {u3011211, u3009618}@connect.hku.hk, dongjingwei@stu.xjtu.edu.cn, u3638070@connect.hku.hk, nlp.lilei@gmail.com, sumiler@connect.hku.hk, jiangjy@link.cuhk.edu.hk, {lpk, cwu}@cs.hku.hk
@@ 1 Introduction
In recent years, large language models (LLMs) (Yang et al., 2024; Dubey et al., 2024; Team et al., 2023; Jiang et al., 2023) have experienced explosive growth in their capabilities, driving significant advancements across various fields (Shao et al., 2023; Guo et al., 2024; Shao et al., 2024). This progress has also sparked interest in developing large vision-language models (LVLMs) (Bai et al., 2025; Chen et al., 2024b; Bai et al., 2023), which, like LLaVA (Li et al., 2024b), have achieved remarkable performance in multi-modal understanding tasks. However, state-of-the-art (SOTA) LVLMs still struggle to handle visual understanding with textual reasoning simultaneously due to inherent modality differences. For example, Ghosh et al. (2024) demonstrate that LVLMs often rely more on their prior language knowledge, neglecting visual information in multi-modal reasoning tasks, such as visual chart understanding and math reasoning, resulting in performance degradation. Figure 2.b illustrates a typical case of this issue, where the reasoning process remains irrelevant to the image.
@@ 4.2 Main Results
ProReason exhibits significant and consistent performance enhancement over baselines across all the benchmarks. As listed in Table 1, despite better performance than the direct method on MME dataset, VDGD and CCoT fail to demonstrate consistent improvements on the other datasets. In contrast, ProReason consistently surpasses all other baselines across all benchmarks for every base model, enhancing the average performance of GPT-4o-mini by 7.9%, demonstrating the superiority and task robustness of ProReason. Section A.5.7 and Table 13 evaluate ProReason on additional benchmarks, including reasoning and VQA tasks, further proving its strong performance and effectiveness across various multi-modal tasks.
@@
Decoupling the visual perception and textual reasoning capabilities of an LVLM outperforms their simultaneous inherent usage. Table 1 illustrates that CoT, utilizing both capabilities simultaneously, does not exhibit consistent performance enhancements over the "Direct" method when applied to the Llama3-LLaVA-NeXT-8B and Qwen2.5-VL-7B-Instruct models. In contrast, despite the same models, ProReason alternates between visual information acquisition and textual reasoning processes, allowing to leverage each capability more effectively. This enables ProReason to consistently outperform CoT with both Llama3-LLaVA-NeXT-8B and GPT-4o-mini across all benchmarks, validating the effectiveness of capability decoupling.
@@
Proactive information acquisition surpasses peer passive methods, especially in complex visual reasoning tasks. Specifically, compared to MME, MathVista and HallusionBench present higher image complexity and question difficulty, and thus require stronger visual understanding and textual reasoning capabilities. This leads to performance degradation of passive methods (i.e., VDGD and CCoT), highlighting their limited applicability to complex visual reasoning tasks. In contrast, ProReason achieves notable performance improvements, up to 5.1%, by proactively acquiring visual information from images rather than generating question-agnostic captions. This aligns with out previous obervations in Section 2 that passive methods introduce substantial information redundancy or omission, misleading subsequent reasoning processes.
@@
Text-only LLMs can be effectively integrated into ProReason for dramatically enhanced performance. As mentioned in Section 3.3, the decoupled visual perception and textual reasoning capabilities facilitate the seamless integration of text-only LLMs. To demonstrate the utility of this advantage, we fix the Vision Expert and replace the textual sub-agents in ProReason with text-only LLMs, according to the configuration in Table 14. As listed in Table 1, with the assistance of powerful existing LLMs, the Llama3-LLaVA-NeXT-8B Vision Expert receives remarkable performance boost across all benchmarks, particularly by 15% on MMMU and 11.7% on MathVista, compared to directly providing answers. In addition, by configuring the Summarizer to Qwen3-32B, ProReason acquires the same kind of deep reasoning capability as Deekseek-R1 and OpenAI o1, resulting in a 10.2% average performance improvement for the Qwen2.5-VL-7B-Instruct Vision Expert. In contrast, ReAct gains a much smaller improvement. This highlights the unique advantage of ProReason in leveraging existing text-only LLMs for enhanced performance. Notably, this advantage may open new avenues for continuously pushing the performance limits of LVLMs with the assistance of existing powerful LLMs.
@@ 4.3 Downstream Task Model Enhancement
High-quality visual reasoning data produced by ProReason significantly improves the performance of downstream task models. We select the test set that has no answer in the MMMU dataset and gather reasoning processes paired with corresponding answers on this set, generated by two configurations detailed in Table 14: ProReason + Qwen72B Assisted and ProReason + Qwen3 Assisted. After eliminating samples with inconsistent answers between these two configurations, we obtain two filtered datasets each containing 5,980 entries. These datasets are then used to fine-tune Qwen2.5-VL-7B-Instruct (Section A.4), yielding two models: ProReason-VL and ProReason-Q3. Notably, ProReason-Q3 inherits the deep reasoning capabilities of Qwen3-32B, which incorporates a reasoning process containing a <think>... </think> part.
@@ 4.4 Efficiency and Complexity Analysis
ProReason improves both efficiency and performance compared to the baseline method. Table 3 presents the evaluation of average token consumption and time expenditure of different methods on the MME and MathVista datasets. Notably, ProReason requires significantly fewer tokens than ReAct, while achieving superior performance over ReAct as analyzed in Section 4.2. Moreover, compared to the visual reasoning frameworks VDGD and CCoT, which also involve multiple image inputs, ProReason’s token consumption on MME is only about 20% higher, yet it achieves a 12.7% performance improvement. Regarding time efficiency, ProReason achieves 11.1% higher performance than VDGD and CCoT on MME with comparable time consumption (18.4s/sample vs. VDGD’s 12.3s and CCoT’s 12.8s), while ReAct is slower than ProReason across datasets, highlighting ProReason’s efficiency advantage. Furthermore, as discussed in Section 4.3, ProReason’s robust visual reasoning capabilities can be transferred to downstream task models, further ensuring the performance and efficiency of our approach.
ProReason achieves a balance between performance and efficiency through its Memory component and adaptive mechanisms. The Memory component enables ProReason to keep compact information representations, avoiding lengthy reasoning traces like ReAct, reducing token use and boosting efficiency. Additionally, ProReason dynamically adjusts Proactive Visual Perception iterations based on question difficulty, minimizing overhead for simple tasks while enhancing complex problem-solving. As shown in Table 12, MathVista’s greater challenge prompts more iterations compared to MME, resulting in higher token usage and longer reasoning times. This aligns with the difficulty levels of the datasets, demonstrating the adaptive nature of ProReason.
@@ A.4 Model Training
[…]
@@ A.5.1 Relative Importance of Sub-agents
Summarizer is the most crucial sub-agent, closely followed by Referee. The replacement of Summarizer results in the most notable performance decline on both MME and MMMU tasks, reaching 6.2% and 10.6%, respectively. This highlights the critical function of the Summarizer in integrating all available information to conclude final answers. Besides, the substitution of Referee leads to a 10.1% reduction on MMMU. Given that MMMU is more challenging than MME, this finding underscores the essential role of the Referee in assessing the sufficiency of information, particularly in more complex visual reasoning tasks. The analysis in Section A.5.5 also demonstrates that the Referee plays a crucial role in enabling ProReason to accurately interpret visual detail information.
Relatively, Dispatcher and Insight Expert are the least essential sub-agents. Specifically, despite a decline, these two sub-agents exhibit significantly less performance degradation than other sub-agents. This can be attributed to the easier task of the Dispatcher, which requires minimal textual reasoning capabilities, and the infrequent calls of the Insight Expert, which is only activated when additional information needs to be inferred—a situation that is rare in current benchmarks. Besides, both sub-agents operate within the acquisition loop, allowing for greater error tolerance. Even if some error occurs, subsequent iterations can compensate for the missing information.
@@
Llama3- LLaVA- NeXT-8B Dispatcher 88.8 (-1.6) 60.9 (-0.7)
@@ A.5.2 Which One is More Crucial: Visual Understanding or Textual Reasoning?
Textual reasoning capabilities outweigh visual understanding for multi-modal reasoning tasks, although both are important. As shown in Table 7, replacing either the Vision Expert or the other agents with the more capable GPT-4o-mini achieves consistent performance enhancement, highlighting the significance of both capabilities. However, substituting the textual sub-agents with GPT-4o-mini results in a more substantial performance boost compared to replacing the Vision Expert. This underscores the greater importance of textual reasoning over visual understanding for multimodal reasoning tasks, aligning with our previous analysis in Section A.5.1 that identifies the Summarizer and Referee as the most crucial sub-agents.
@@
GPT-4o -mini Textual Sub-Agents 84.7 (+13.2) 54.5 (+2.0)
@@
Llama3- LLaVA- NeXT-8B All Sub-Agents 71.5 52.5
@@ A.5.3 The Critical Implication of Decoupling
Decoupling serves as a crucial mechanism for improving ProReason’s performance in complex visual reasoning tasks. As shown in Table 8, merging Vision and Insight Experts results in a 3% performance drop on the MMMU benchmark, while combining Proactive Visual Perception with Textual Reasoning leads to a more significant 4.8% reduction (56.8 vs. CoT’s 58.5). Although agent merging also causes performance declines on MME, the merged versions still outperform CoT. Notably, given MMMU’s substantially higher complexity compared to MME, these findings reveal that the decoupling of visual-textual processing fundamentally drives its performance gains in complex scenarios.
@@ A.5.5 Referee’s Dispel of Hallucinations
The Referee module effectively filters hallucinated information to enhance the visual comprehension capabilities of our framework. As demonstrated in Table 10, the performance improvement on HallusionBench (2.8%) significantly outpaces that on MMMU (1.5%) as attempt opportunities increase from 1 to 5. Given HallusionBench’s dual emphasis on reasoning proficiency and precise evaluation of visual details/hallucination control compared to MMMU, these results suggest that the Referee mechanism can effectively identify erroneous or irrelevant visual information, thereby strengthening ProReason’s capacity for meticulous visual understanding.

### v5

@@ Abstract
Large vision-language models (LVLMs) have witnessed significant progress on visual understanding tasks. However, they often prioritize language knowledge over image information on visual reasoning tasks, incurring performance degradation. To tackle this issue, we first identify the drawbacks of existing solutions (i.e., limited multi-modal reasoning capacities, and insufficient and irrelevant visual descriptions). We then decompose visual reasoning process into two stages: proactive visual perception (i.e., eyesight) and textual reasoning (i.e., wisdom), and introduce a novel visual reasoning framework named ProReason. This framework features decoupled vision-reasoning capabilities and multi-run proactive perception. Briefly, given a multi-modal question, ProReason iterates proactive information collection and reasoning until the answer can be concluded with necessary and sufficient visual descriptions. Notably, the disassociation of capabilities allows seamless integration of existing large language models (LLMs) to compensate for the reasoning deficits of LVLMs. Our extensive experiments demonstrate that ProReason outperforms existing multi-step reasoning frameworks on various benchmarks for both open-source and closed-source models, with the average performance gain reaching 13.2%. Besides, the integration of LLMs allows ProReason to produce high-quality visual reasoning data, which empowers ProReason-distilled models (i.e., ProReason-VL and ProReason-Q3) to achieve superior performance in downstream tasks. Our insights into existing solutions and the decoupled perspective for feasible integration of LLMs illuminate future research on visual reasoning techniques, especially LLM-assisted ones. The code is available at https://github.com/lian-tian-mo-zun/Pro_Reason.
@@
Jingqi Zhou, Sheng Wang, Jingwei Dong, Kai Liu, Lei Li, Jiahui Gao, Jiyue Jiang, Lingpeng Kong, Chuan Wu The University of Hong Kong, The Chinese University of Hong Kong, {u3011211, u3009618, u3013005, u3638070}@connect.hku.hk, nlp.lilei@gmail.com, sumiler@connect.hku.hk, jiangjy@link.cuhk.edu.hk, {lpk, cwu}@cs.hku.hk
@@ 1 Introduction
In recent years, large language models (LLMs) (Yang et al., 2024; Dubey et al., 2024; Team et al., 2023; Jiang et al., 2023) have experienced explosive growth in their capabilities, driving significant advancements across various fields (Shao et al., 2023; Guo et al., 2024; Shao et al., 2024). This progress has also sparked interest in developing large vision-language models (LVLMs) (Xiaomi and Team, 2025; Bai et al., 2025; Chen et al., 2024b; Bai et al., 2023), which, like LLaVA (Li et al., 2024b), have achieved remarkable performance in multi-modal understanding tasks. However, state-of-the-art (SOTA) LVLMs still struggle to handle visual understanding with textual reasoning simultaneously due to inherent modality differences. For example, Ghosh et al. (2024) demonstrate that LVLMs often rely more on their prior language knowledge, neglecting visual information in multi-modal reasoning tasks, such as visual chart understanding and math reasoning, resulting in performance degradation. Figure 2.b illustrates a typical case of this issue, where the reasoning process remains irrelevant to the image.
@@ 4.2 Main Results
ProReason exhibits significant and consistent performance enhancement over baselines across all the benchmarks.
@@
As listed in Table 1, despite better performance than the direct method on MME dataset, VDGD and CCoT fail to demonstrate consistent improvements on the other datasets. In contrast, ProReason consistently surpasses all other baselines across all benchmarks for every base model, enhancing the average performance of GPT-4o-mini by 7.9%, demonstrating the superiority and task robustness of ProReason. Section A.5.7 and Table 13 evaluate ProReason on additional benchmarks, including reasoning and VQA tasks, further proving its strong performance and effectiveness across various multi-modal tasks.
Decoupling the visual perception and textual reasoning capabilities of an LVLM outperforms their simultaneous inherent usage.
@@
Table 1 illustrates that CoT, utilizing both capabilities simultaneously, does not exhibit consistent performance enhancements over the "Direct" method when applied to the Llama3-LLaVA-NeXT-8B and Qwen2.5-VL-7B-Instruct models. In contrast, despite the same models, ProReason alternates between visual information acquisition and textual reasoning processes, allowing to leverage each capability more effectively. This enables ProReason to consistently outperform CoT with both Llama3-LLaVA-NeXT-8B and GPT-4o-mini across all benchmarks, validating the effectiveness of capability decoupling.
Proactive information acquisition surpasses peer passive methods
@@
, especially in complex visual reasoning tasks. Specifically, compared to MME, MathVista and HallusionBench present higher image complexity and question difficulty, and thus require stronger visual understanding and textual reasoning capabilities. This leads to performance degradation of passive methods (i.e., VDGD and CCoT), highlighting their limited applicability to complex visual reasoning tasks. In contrast, ProReason achieves notable performance improvements, up to 5.1%, by proactively acquiring visual information from images rather than generating question-agnostic captions. This aligns with out previous obervations in Section 2 that passive methods introduce substantial information redundancy or omission, misleading subsequent reasoning processes.
Text-only LLMs can be effectively integrated into ProReason for dramatically enhanced performance.
@@
As mentioned in Section 3.3, the decoupled visual perception and textual reasoning capabilities facilitate the seamless integration of text-only LLMs. To demonstrate the utility of this advantage, we fix the Vision Expert and replace the textual sub-agents in ProReason with text-only LLMs, according to the configuration in Table 15. As listed in Table 1, with the assistance of powerful existing LLMs, the Llama3-LLaVA-NeXT-8B Vision Expert receives remarkable performance boost across all benchmarks, particularly by 15% on MMMU and 11.7% on MathVista, compared to directly providing answers. In addition, by configuring the Summarizer to Qwen3-32B, ProReason acquires the same kind of deep reasoning capability as Deekseek-R1 and OpenAI o1, resulting in a 10.2% average performance improvement for the Qwen2.5-VL-7B-Instruct Vision Expert. In contrast, ReAct gains a much smaller improvement. This highlights the unique advantage of ProReason in leveraging existing text-only LLMs for enhanced performance. Notably, this advantage may open new avenues for continuously pushing the performance limits of LVLMs with the assistance of existing powerful LLMs.
@@ 4.3 Downstream Task Model Enhancement
High-quality visual reasoning data produced by ProReason significantly improves the performance of downstream task models.
We select the test set that has no answer in the MMMU dataset and gather reasoning processes paired with corresponding answers on this set, generated by two configurations detailed in Table 15: ProReason + Qwen72B Assisted and ProReason + Qwen3 Assisted. After eliminating samples with inconsistent answers between these two configurations, we obtain two filtered datasets each containing 5,980 entries. These datasets are then used to fine-tune Qwen2.5-VL-7B-Instruct (Section A.4), yielding two models: ProReason-VL and ProReason-Q3. Notably, ProReason-Q3 inherits the deep reasoning capabilities of Qwen3-32B, which incorporates a reasoning process containing a <think>... </think> part.
@@ 4.4 Efficiency and Complexity Analysis
ProReason improves both efficiency and performance compared to the baseline method
. Table 3 presents the evaluation of average token consumption and time expenditure of different methods on the MME and MathVista datasets. Notably, ProReason requires significantly fewer tokens than ReAct, while achieving superior performance over ReAct as analyzed in Section 4.2. Moreover, compared to the visual reasoning frameworks VDGD and CCoT, which also involve multiple image inputs, ProReason’s token consumption on MME is only about 20% higher, yet it achieves a 12.7% performance improvement. Regarding time efficiency, ProReason achieves 11.1% higher performance than VDGD and CCoT on MME with comparable time consumption (18.4s/sample vs. VDGD’s 12.3s and CCoT’s 12.8s), while ReAct is slower than ProReason across datasets, highlighting ProReason’s efficiency advantage. Furthermore, as discussed in Section 4.3, ProReason’s robust visual reasoning capabilities can be transferred to downstream task models, further ensuring the performance and efficiency of our approach.
ProReason achieves a balance between performance and efficiency through its Memory component and adaptive mechanisms
. The Memory component enables ProReason to keep compact information representations, avoiding lengthy reasoning traces like ReAct, reducing token use and boosting efficiency. Additionally, ProReason dynamically adjusts Proactive Visual Perception iterations based on question difficulty, minimizing overhead for simple tasks while enhancing complex problem-solving. As shown in Table 12, MathVista’s greater challenge prompts more iterations compared to MME, resulting in higher token usage and longer reasoning times. This aligns with the difficulty levels of the datasets, demonstrating the adaptive nature of ProReason.
@@ 7 Ethics Statement
8 Acknowledgements
We want to thank our anonymous AC and reviewers for their feedback. This work was supported in part by Hong Kong Innovation and Technology Commission’s Innovation and Technology Fund (Award No. ITS/269/22FP), Hong Kong RGC grants C7004-22G (CRF) and CRS_PolyU501/23 (CRS).
@@ A.4 Model Training
we use Supervised Fine-Tuning(SFT) and employ the parameter-efficient fine-tuning method LoRA (Hu et al., 2021). Specifically, we uniformly set the learning rate to $1\times 10^{-4}$, $\texttt{lora\_dropout}=0$, and train the ProReason-VL for 1 epoch and ProReason-Q3 for 2 epoch. These parameters are the optimal values obtained after multiple attempts.
@@ A.5.1 Relative Importance of Sub-agents
Summarizer is the most crucial sub-agent, closely followed by Referee.
The replacement of Summarizer results in the most notable performance decline on both MME and MMMU tasks, reaching 6.2% and 10.6%, respectively. This highlights the critical function of the Summarizer in integrating all available information to conclude final answers. Besides, the substitution of Referee leads to a 10.1% reduction on MMMU. Given that MMMU is more challenging than MME, this finding underscores the essential role of the Referee in assessing the sufficiency of information, particularly in more complex visual reasoning tasks. The analysis in Section A.5.5 also demonstrates that the Referee plays a crucial role in enabling ProReason to accurately interpret visual detail information.
Relatively, Dispatcher and Insight Expert are the least essential sub-agents.
Specifically, despite a decline, these two sub-agents exhibit significantly less performance degradation than other sub-agents. This can be attributed to the easier task of the Dispatcher, which requires minimal textual reasoning capabilities, and the infrequent calls of the Insight Expert, which is only activated when additional information needs to be inferred—a situation that is rare in current benchmarks. Besides, both sub-agents operate within the acquisition loop, allowing for greater error tolerance. Even if some error occurs, subsequent iterations can compensate for the missing information.
@@
Llama3-
LLaVA-
NeXT-8B
Dispatcher 88.8 (-1.6) 60.9 (-0.7)
@@ A.5.2 Which One is More Crucial: Visual Understanding or Textual Reasoning?
Textual reasoning capabilities outweigh visual understanding for multi-modal reasoning tasks, although both are important.
As shown in Table 7, replacing either the Vision Expert or the other agents with the more capable GPT-4o-mini achieves consistent performance enhancement, highlighting the significance of both capabilities. However, substituting the textual sub-agents with GPT-4o-mini results in a more substantial performance boost compared to replacing the Vision Expert. This underscores the greater importance of textual reasoning over visual understanding for multimodal reasoning tasks, aligning with our previous analysis in Section A.5.1 that identifies the Summarizer and Referee as the most crucial sub-agents.
@@
GPT-4o
-mini
Textual Sub-Agents 84.7 (+13.2) 54.5 (+2.0)
@@
Llama3-
LLaVA-
NeXT-8B
All Sub-Agents 71.5 52.5
@@ A.5.3 The Critical Implication of Decoupling
Decoupling serves as a crucial mechanism for improving ProReason’s performance in complex visual reasoning tasks
. As shown in Table 8, merging Vision and Insight Experts results in a 3% performance drop on the MMMU benchmark, while combining Proactive Visual Perception with Textual Reasoning leads to a more significant 4.8% reduction (56.8 vs. CoT’s 58.5). Although agent merging also causes performance declines on MME, the merged versions still outperform CoT. Notably, given MMMU’s substantially higher complexity compared to MME, these findings reveal that the decoupling of visual-textual processing fundamentally drives its performance gains in complex scenarios.
@@ A.5.5 Referee’s Dispel of Hallucinations
The Referee module effectively filters hallucinated information to enhance the visual comprehension capabilities of our framework
. As demonstrated in Table 10, the performance improvement on HallusionBench (2.8%) significantly outpaces that on MMMU (1.5%) as attempt opportunities increase from 1 to 5. Given HallusionBench’s dual emphasis on reasoning proficiency and precise evaluation of visual details/hallucination control compared to MMMU, these results suggest that the Referee mechanism can effectively identify erroneous or irrelevant visual information, thereby strengthening ProReason’s capacity for meticulous visual understanding.
@@ A.5.7 Performance Evaluation of ProReason on Supplementary Datasets
A.5.8 The Impact of Iterative Reasoning on ProReason’s Performance
As shown in the Table 14, we evaluated the performance of ProReason, powered by GPT-4o-mini, on the MME, MMMU, and MathVista datasets, given different maximum loop iterations. The performance of ProReason is observed to enhance, as the maximum iterations increase up to three, after which the performance tends to stabilize. This observation indicates that most questions in the MME, MMMU, and MathVista benchmarks can be effectively addressed within three cycles of question-oriented visual information extraction. In cases of a few difficult questions, ProReason can adaptively extend the number of iterations, proactively gathering more information from the image to arrive at the solution.